OpenStudy (anonymous):

\[Find the derivative of the function f(x)=3e^{x} - x ^{4}\]

OpenStudy (abb0t):

\[\frac{ d }{ dx }e^x = e^x\] \[\frac{ d }{ dx } a^n = na^{n-1}\]

OpenStudy (anonymous):

so, f'(x)=\[3e^{x} - 4x ^{3}\]

OpenStudy (abb0t):

Yes.

OpenStudy (anonymous):

thanks